# Relatively big projective modules and their applications to direct sum decompositions

### v1

@@ 6.2. Locally semiperfect algebras over $h$-local domains
Note that if $U$ is a projective $\Lambda_{\mathfrak{m}}$-module and $I_{\mathfrak{m}}\subseteq\operatorname{Tr}_{\Lambda}(U)$, then $U$ generates any $\Lambda_{\mathfrak{m}}$-module $W$ satisfying $WI_{\mathfrak{m}}=W$. If $W$ is also finitely generated, then there exists $k\in\mathbb{N}$ such that $W$ is a homomorphic image of $U^{k}$. Therefore $V_{\mathfrak{m}}$ is a homomorphic image of $\bigoplus_{i\in C}U_{i}$ whenever $C$ is a cofinal subset of $\mathbb{N}$. In particular, there exists a homomorphism $g\in\operatorname{Hom}_{\Lambda_{\mathfrak{m}}}(P_{\mathfrak{m}},V_{\mathfrak{m}})$ which is onto and $g(P_{1})=0$ (recall $P$ is embedded in $P_{\mathfrak{m}}$). Since $V$ is finitely generated, there exists $P_{0}^{\prime}$, a finitely generated submodule of $P_{\mathfrak{m}}$, such that $g(P_{0}^{\prime})=V_{\mathfrak{m}}$. Scaling by a suitable element of $R\setminus\mathfrak{m}$ (which is invertible over $\Lambda_{\mathfrak{m}}$) we observe that there exists finitely generated $P_{0}\subseteq P$ such that $P_{0}\Lambda_{\mathfrak{m}}=P_{0}^{\prime}$. Let $F$ be a finite subset of $\omega$ such that $\nu(P_{0}+P_{1})$ is contained in $\Lambda^{(F)}$, further let $\iota\in\operatorname{Hom}_{\Lambda}(\Lambda^{(F)},\Lambda^{(\omega)})$ and $\varrho\in\operatorname{Hom}_{\Lambda}(\Lambda^{(\omega)},\Lambda^{(F)})$ be the canonical embedding and projection. Since $\Lambda^{(F)}$ is finitely generated over $\Lambda$ and torsion-free over $R$, by Lemma 5.2 there exists $h\in\operatorname{Hom}_{\Lambda}(\Lambda^{(F)},V)$ and $q\in R_{\mathfrak{m}}$ such that $h\otimes q=g(\pi\iota)_{\mathfrak{m}}$. Consider $f:=h\varrho\nu$. Then $f\otimes q=h\otimes q\circ(\varrho\nu)\otimes 1=g(\pi\iota\varrho\nu)_{\mathfrak{m}}$. Since $\Lambda^{(F)}$ contains $\nu(P_{0}+P_{1})$, $\pi\iota\varrho\nu$ is identity on $P_{0}+P_{1}$ and, consequently, $f\otimes q$ is onto and $f\otimes q(P_{1}\otimes 1)=0$. This implies that $f_{\mathfrak{m}}$ is onto and $f(P_{1})=0$.
@@
We are going to apply Lemma 6.12 in the following situation: $\Lambda$ is a locally semiperfect torsion-free algebra over an $h$-local domain $R$ such that $\Lambda_{Q}$ is a simple artinian ring. Assume that $P$ is a countably generated projective $\Lambda$-module and $0\neq I$ is the least element of the set of ideals $\{J\unlhd\Lambda\mid P/PJ\text{ is finitely generated}\}$. Take any $\mathfrak{m}\in\operatorname{mSpec}(R)$ and let us check that $P_{\mathfrak{m}}$ is a countable direct sum of finitely generated projective $\Lambda_{\mathfrak{m}}$-modules such that almost all these direct summands have trace equal to $I_{\mathfrak{m}}$.
Indeed, since $P$ is not finitely generated, $P_{\mathfrak{m}}$ is not a finitely generated $\Lambda_{\mathfrak{m}}$-module by Proposition 6.11. Since $\Lambda_{\mathfrak{m}}$ is semiperfect, $P_{\mathfrak{m}}$ is a direct sum of finitely generated modules, $P_{\mathfrak{m}}=U_{1}^{m_{1}}\oplus\cdots\oplus U_{t}^{m_{t}}\oplus(U_{t+1}\oplus\cdots\oplus U_{s})^{(\omega)}$, where $U_{1},\dots,U_{s}$ is a representative set of indecomposable projective $\Lambda_{\mathfrak{m}}$-modules. Let $U^{\prime}=U_{t+1}\oplus\cdots\oplus U_{s}$. As explained in the proof of Proposition 6.11(2), $I_{\mathfrak{m}}=\operatorname{Tr}_{\Lambda_{\mathfrak{m}}}(U^{\prime})$. Therefore $P_{\mathfrak{m}}=U_{1}^{m_{1}}\oplus\cdots\oplus U_{t}^{m_{t}}\oplus U^{\prime(\omega)}$ is a decomposition of $P_{\mathfrak{m}}$ having the required properties.
Lemma 6.12 can be stated in simpler form when $\mathfrak{m}=(0)$ and $I=\Lambda$.
@@
Let $R$ be a commutative domain with field of fractions $Q$, and let $\Lambda$ be a torsion-free $R$-algebra. Let $P$ be a projective $\Lambda$-module such that $P_{Q}$ is countably generated but not finitely generated free, and $V$ a finitely generated $\Lambda$-module torsion-free over $R$. Then for any $P_{1}$ finitely generated submodule of $P$, there exists $f\in\operatorname{Hom}_{\Lambda}(P,V)$ such that $f_{Q}$ is onto and $f(P_{1})=0$.
@@
Let $R$ be an $h$-local domain with field of fractions $Q$, and let $\Lambda$ be a locally semiperfect torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian. Let $P$ be a countably generated projective $\Lambda$-module such that $I=\min\{J\unlhd\Lambda\mid P/PJ\text{ is finitely generated}\}$ exists and it is the trace of a countably generated projective module. Then $P$ is $I$-big.
Proof.  We have to prove that if $U$ is a countably generated projective right $\Lambda$-module and $\operatorname{Tr}_{\Lambda}(U)\subseteq I$, then $U$ is a direct summand of $P$. By Corollary 6.7, $U^{(\omega)}$ is a direct sum of finitely generated modules, so we may assume that $U$ is a direct sum of finitely generated modules.
@@
Assume we have proved the claim and let $U=\bigoplus_{i\in\mathbb{N}}U_{i}$ with $U_{i}$ finitely generated. Let $p_{1},p_{2},\dots\in P$ be a countable set of generators of $P$. Consider the following construction: Let $f_{1}\colon P\to U_{1}$ be an epimorphism and let $P_{1}$ be a finitely generated submodule of $P$ such that $f_{1}(P_{1})=U_{1}$ and $p_{1}\in P_{1}$. Assume that $f_{1},f_{2},\dots,f_{k}$ and $P_{1},P_{2},\dots,P_{k}$ were defined such that $f_{i}\colon P\to U_{i}$ is an epimorphism, $P_{1}\subseteq\cdots\subseteq P_{k}$ are finitely generated submodules of $P$ such that $f_{i}(P_{i-1})=0$, $p_{i}\in P_{i}$, and $f_{i}(P_{i})=U_{i}$ (we can set $P_{0}=0$). By the claim, there exists an epimorphism $f_{k+1}\colon P\to U_{k+1}$ such that $f_{k+1}(P_{k})=0$. Finally, let $P_{k+1}$ be a finitely generated submodule of $P$ satisfying $P_{k}\subseteq P_{k+1}$, $p_{k+1}\in P_{k+1}$ and $f_{k+1}(P_{k+1})=U_{k+1}$.
Consider the canonical homomorphism $f=\prod_{k\in\mathbb{N}}f_{k}\colon P\to\prod_{k}U_{k}$. Observe that $f_{k}(p_{l})=0$ if $k>l$ so ${\rm Im}\ f\subseteq\bigoplus_{i\in\mathbb{N}}U_{i}$. An easy inductive argument also gives $f(P_{i})=\bigoplus_{j=1}^{i}U_{j}$ for any $i\in\mathbb{N}$. Thus $f$ is an epimorphism.
Thus we are left to prove the claim. First note that we may assume $I\neq 0$ since any countably generated projective module is 00-big. Then $P$ is not finitely generated and therefore, by Proposition 6.11, $P_{Q}$ cannot be finitely generated. Hence $P_{Q}\simeq\Lambda_{Q}^{(\omega)}$.
By Lemma 6.14, there exists $g\in\operatorname{Hom}_{\Lambda}(P,V)$ such that $g_{Q}$ is onto and $g(P_{0})=0$. Let $T_{0}$ be a finitely generated submodule of $P$ containing $P_{0}$ such that $g_{Q}(T_{0}\Lambda_{Q})=(g(T_{0}))_{Q}=V_{Q}$. Since $V$ is finitely generated, $V_{\mathfrak{m}}=g(T_{0})_{\mathfrak{m}}$ for almost all maximal ideals $\mathfrak{m}$ of $R$. Let $\mathfrak{m}_{1},\mathfrak{m}_{2},\dots,\mathfrak{m}_{k}$ be the maximal ideals of $R$ where the equality does not hold. If there are no such ideals then we are done, since $g$ is onto. So we may assume that $k\geq 1$.
@@
Then $f=g+\sum_{i=1}^{k}f_{i}\in\operatorname{Hom}_{\Lambda}(P,V)$ satisfies $f(T_{0})=g(T_{0})$ hence $f_{\mathfrak{m}}$ is onto if $\mathfrak{m}$ is not among $\mathfrak{m}_{1},\dots,\mathfrak{m}_{k}$. Moreover, $f(T_{i})=(g+\sum_{j=1}^{i}f_{j})(T_{i})$ shows that $f_{\mathfrak{m}_{i}}$ is onto for every $i=1,\dots,k$. This implies that $f$ is onto.
Assume $f_{1},\dots,f_{\ell-1}$ and $T_{0},T_{1},\dots,T_{\ell-1}$ are found. Consider $g_{\ell}=g+\sum_{i=1}^{\ell-1}f_{i}$. Recall that $V_{\mathfrak{m}_{\ell}}$ is a finitely generated projective module over a semiperfect ring $\Lambda_{\mathfrak{m}_{\ell}}$. Then $V_{\mathfrak{m}_{\ell}}/V_{\mathfrak{m}_{\ell}}J(\Lambda_{\mathfrak{m}_{\ell}})$ is a finitely generated (semisimple module), so if $\pi\colon V_{\mathfrak{m}_{\ell}}\to V_{\mathfrak{m}_{\ell}}/V_{\mathfrak{m}_{\ell}}J(\Lambda_{\mathfrak{m}_{\ell}})$ is the canonical projection, $\pi((g_{\ell}(P))_{\mathfrak{m}_{\ell}})$ is a finitely generated submodule of $V_{\mathfrak{m}_{\ell}}/V_{\mathfrak{m}_{\ell}}J(\Lambda_{\mathfrak{m}_{\ell}})$. Let $T_{\ell}^{\prime}\subseteq P$ be a finitely generated submodule of $P$ containing $T_{\ell-1}$ such that $(g_{\ell}(T_{\ell}^{\prime}))_{\mathfrak{m}_{\ell}}+V_{\mathfrak{m}_{\ell}}J(\Lambda_{\mathfrak{m}_{\ell}})=(g_{\ell}(P))_{\mathfrak{m}_{\ell}}+V_{\mathfrak{m}_{\ell}}J(\Lambda_{\mathfrak{m}_{\ell}})$.
By Lemma 6.12, there exists $f_{\ell}\in{\rm Hom}_{\Lambda}(P,V)$ such that $(f_{\ell})_{\mathfrak{m}_{\ell}}$ is onto and $f_{\ell}(T_{\ell}^{\prime})=0$. Then the image of $g_{\ell}+f_{\ell}$ contains $g_{\ell}(T_{\ell}^{\prime})$ hence also $\mbox{\rm Im}\ (g_{\ell}+f_{\ell})_{\mathfrak{m}_{\ell}}+V_{\mathfrak{m}_{\ell}}J(\Lambda_{\mathfrak{m}_{\ell}})$ contains $\mbox{\rm Im}\ (g_{\ell})_{\mathfrak{m}_{\ell}}$ and since $(f_{\ell})_{\mathfrak{m}_{\ell}}$ is onto, $\mbox{\rm Im}\ (g_{\ell}+f_{\ell})_{\mathfrak{m}_{\ell}}+V_{\mathfrak{m}_{\ell}}J(\Lambda_{\mathfrak{m}_{\ell}})=V_{\mathfrak{m}_{\ell}}$. Nakayama’s Lemma gives that $(g_{\ell}+f_{\ell})_{\mathfrak{m}_{\ell}}$ is onto and therefore there exists $T_{\ell}$ finitely generated submodule of $P$ containing $T_{\ell}^{\prime}$ such that $((g_{\ell}+f_{\ell}))_{\mathfrak{m}_{\ell}}((T_{\ell})_{\mathfrak{m}_{\ell}})=V_{\mathfrak{m}_{\ell}}$, that is $(iii)$ holds if $i=\ell$.
@@
Proof.  Propositions 6.11 and 6.15 show that $V_{r}^{*}(\Lambda)=V_{r}(\Lambda)\sqcup B_{r}(\Lambda)$. The symmetry of the hypothesis implies that $V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)=V_{\ell}^{*}(\Lambda)$. By Remark 6.9, we can deduce from Corollary 2.12 that both monoids are isomorphic.
Hinohara [hinohara] proved that every projective module over an $h$-local (or more generally, weakly noetherian) domain is either finitely generated or free. We can extend this result to our setting.
@@
[hinohara] Let $R$ be an $h$-local domain with field of fractions $Q$, and let $\Lambda$ be a torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian. For each maximal ideal $\mathfrak{m}$ of $R$, assume that $\Lambda_{\mathfrak{m}}$ is Morita equivalent to a local ring. Then any infinitely generated projective $\Lambda$-module that is not finitely generated is free.
Proof.  By Remark 6.9, ${\mathcal{T}}(\Lambda)=\{0,\Lambda\}$, and hence $V(\Lambda)\sqcup B(\Lambda)=V(\Lambda)\sqcup\{\Lambda^{(\omega)}\}$. The result then follows from Theorem 6.16.

### v2

@@ 6.2. Locally semiperfect algebras over $h$-local domains
Note that if $U$ is a projective $\Lambda_{\mathfrak{p}}$-module and $I_{\mathfrak{p}}\subseteq\operatorname{Tr}_{\Lambda}(U)$, then $U$ generates any $\Lambda_{\mathfrak{p}}$-module $W$ satisfying $WI_{\mathfrak{p}}=W$. If $W$ is also finitely generated, then there exists $k\in\mathbb{N}$ such that $W$ is a homomorphic image of $U^{k}$. Therefore $V_{\mathfrak{p}}$ is a homomorphic image of $\bigoplus_{i\in C}U_{i}$ whenever $C$ is a cofinal subset of $\mathbb{N}$. In particular, there exists an onto homomorphism $g\in\operatorname{Hom}_{\Lambda_{\mathfrak{p}}}(P_{\mathfrak{p}},V_{\mathfrak{p}})$ such that $g(P_{1})=0$ (recall that $P$ is embedded in $P_{\mathfrak{p}}$). Since $V$ is finitely generated, there exists $P_{0}^{\prime}$, a finitely generated submodule of $P_{\mathfrak{p}}$, such that $g(P_{0}^{\prime})=V_{\mathfrak{p}}$. Scaling by a suitable element of $R\setminus\mathfrak{p}$ (which is invertible over $\Lambda_{\mathfrak{p}}$), we observe that there exists a finitely generated $P_{0}\subseteq P$ such that $P_{0}\Lambda_{\mathfrak{p}}=P_{0}^{\prime}$.
Let $F$ be a finite subset of $\omega$ such that $\nu(P_{0}+P_{1})$ is contained in $\Lambda^{(F)}$. Further, let $\iota\in\operatorname{Hom}_{\Lambda}(\Lambda^{(F)},\Lambda^{(\omega)})$ and $\varrho\in\operatorname{Hom}_{\Lambda}(\Lambda^{(\omega)},\Lambda^{(F)})$ be the canonical embedding and projection. Since $\Lambda^{(F)}$ is finitely generated over $\Lambda$ and torsion-free over $R$, by Lemma 5.2, there exist $h\in\operatorname{Hom}_{\Lambda}(\Lambda^{(F)},V)$ and $q\in R_{\mathfrak{p}}$ such that $h\otimes q=g(\pi\iota)_{\mathfrak{p}}$. Consider $f:=h\varrho\nu$. Then $f\otimes q=h\otimes q\circ(\varrho\nu)\otimes 1=g(\pi\iota\varrho\nu)_{\mathfrak{p}}$. Since $\Lambda^{(F)}$ contains $\nu(P_{0}+P_{1})$, $\pi\iota\varrho\nu$ is the identity on $P_{0}+P_{1}$, and consequently, $f\otimes q$ is onto and $f\otimes q(P_{1}\otimes 1)=0$. This implies that $f_{\mathfrak{p}}$ is onto and $f(P_{1})=0$.
@@
We are going to apply Lemma 6.13 in the following situation: $\Lambda$ is a locally semiperfect torsion-free algebra over an $h$-local domain $R$ such that $\Lambda_{Q}$ is a simple artinian ring. Assume that $P$ is a countably generated projective $\Lambda$-module and $0\neq I$ is the least element of $\mathcal{I}(P)=\{J\unlhd\Lambda\mid P/PJ\text{ is finitely generated}\}$. Take any $\mathfrak{m}\in\operatorname{mSpec}(R)$ and let us check that $P_{\mathfrak{m}}$ is a countable direct sum of finitely generated projective $\Lambda_{\mathfrak{m}}$-modules such that almost all these direct summands have a trace equal to $I_{\mathfrak{m}}$.
Indeed, since $P$ is not finitely generated, $P_{\mathfrak{m}}$ is not a finitely generated $\Lambda_{\mathfrak{m}}$-module by Proposition 6.12. Since $\Lambda_{\mathfrak{m}}$ is semiperfect, $P_{\mathfrak{m}}$ is a direct sum of finitely generated modules $P_{\mathfrak{m}}=U_{1}^{m_{1}}\oplus\cdots\oplus U_{t}^{m_{t}}\oplus(U_{t+1}\oplus\cdots\oplus U_{s})^{(\omega)}$, where $U_{1},\dots,U_{s}$ is a representative set of indecomposable projective $\Lambda_{\mathfrak{m}}$-modules. Let $U^{\prime}=U_{t+1}\oplus\cdots\oplus U_{s}$. As explained in the proof of Proposition 6.12(ii), $I_{\mathfrak{m}}=\operatorname{Tr}_{\Lambda_{\mathfrak{m}}}(U^{\prime})$. Therefore, $P_{\mathfrak{m}}=U_{1}^{m_{1}}\oplus\cdots\oplus U_{t}^{m_{t}}\oplus U^{\prime(\omega)}$ is a decomposition of $P_{\mathfrak{m}}$ that has the required properties.
Lemma 6.13 can be stated in simpler form when $\mathfrak{p}=(0)$ and $I=\Lambda$.
@@
Let $R$ be a commutative domain with field of fractions $Q$, and let $\Lambda$ be a torsion-free $R$-algebra. Let $P$ be a projective $\Lambda$-module such that $P_{Q}$ is countably generated but not finitely generated free, and let $V$ be a finitely generated $\Lambda$-module which is torsion-free over $R$. Then, for any finitely generated submodule $P_{1}$ of $P$, there exists $f\in\operatorname{Hom}_{\Lambda}(P,V)$ such that $f(P_{1})=0$, and $f_{Q}$ is onto.
@@
Let $R$ be an $h$-local domain with field of fractions $Q$, and let $\Lambda$ be a locally semiperfect torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian. Let $P$ be a countably generated projective $\Lambda$-module such that $I=\min\{J\unlhd\Lambda\mid P/PJ\text{ is finitely generated}\}$ exists and is the trace of a countably generated projective module. Then $P$ is $I$-big.
Proof. We have to prove that if $U$ is a countably generated projective right $\Lambda$-module and $\operatorname{Tr}_{\Lambda}(U)\subseteq I$, then $U$ is a direct summand of $P$. By Corollary 6.8, $U^{(\omega)}$ is a direct sum of finitely generated modules. Thus, we may assume that $U$ is a direct sum of finitely generated modules.
@@
Assume we have proved the claim and let $U=\bigoplus_{i\in\mathbb{N}}U_{i}$, with each $U_{i}$ being finitely generated. Let $p_{1},p_{2},\dots\in P$ be a countable set of generators of $P$. Consider the following construction: Let $f_{1}\colon P\to U_{1}$ be an epimorphism, and let $P_{1}$ be a finitely generated submodule of $P$ such that $f_{1}(P_{1})=U_{1}$ and $p_{1}\in P_{1}$. Assume that $f_{1},f_{2},\dots,f_{k}$ and $P_{1},P_{2},\dots,P_{k}$ are defined such that $f_{i}\colon P\to U_{i}$ is an epimorphism, $P_{1}\subseteq\cdots\subseteq P_{k}$ are finitely generated submodules of $P$ such that $f_{i}(P_{i-1})=0$, $p_{i}\in P_{i}$, and $f_{i}(P_{i})=U_{i}$ (we can set $P_{0}=0$). By the claim, there exists an epimorphism $f_{k+1}\colon P\to U_{k+1}$ such that $f_{k+1}(P_{k})=0$. Finally, let $P_{k+1}$ be a finitely generated submodule of $P$ satisfying $P_{k}\subseteq P_{k+1}$, $p_{k+1}\in P_{k+1}$, and $f_{k+1}(P_{k+1})=U_{k+1}$.
Consider the canonical homomorphism $f=\prod_{k\in\mathbb{N}}f_{k}\colon P\to\prod_{k}U_{k}$. Observe that $f_{k}(p_{l})=0$ if $k>l$, so ${\rm Im}\ f\subseteq\bigoplus_{i\in\mathbb{N}}U_{i}$. An easy inductive argument also gives $f(P_{i})=\bigoplus_{j=1}^{i}U_{j}$ for any $i\in\mathbb{N}$. Thus $f$ is an epimorphism.
Thus, we are left to prove the claim. First, note that we may assume $I\neq 0$ since any countably generated projective module is 0-big. Then $P$ is not finitely generated. Therefore, by Proposition 6.12, $P_{Q}$ cannot be finitely generated. Hence $P_{Q}\simeq\Lambda_{Q}^{(\omega)}$.
By Lemma 6.15, there exists $g\in\operatorname{Hom}_{\Lambda}(P,V)$ such that $g_{Q}$ is onto and $g(P_{0})=0$. Let $T_{0}$ be a finitely generated submodule of $P$ containing $P_{0}$ such that $g_{Q}(T_{0}\Lambda_{Q})=(g(T_{0}))_{Q}=V_{Q}$. Since $V$ is finitely generated, $V_{\mathfrak{m}}=g(T_{0})_{\mathfrak{m}}$ for almost all maximal ideals $\mathfrak{m}$ of $R$. Let $\mathfrak{m}_{1},\mathfrak{m}_{2},\dots,\mathfrak{m}_{k}$ be the maximal ideals of $R$ where the equality does not hold. If there are no such ideals, then we are done, since $g$ is onto. So we may assume that $k\geq 1$.
@@
Then $f=g+\sum_{i=1}^{k}f_{i}\in\operatorname{Hom}_{\Lambda}(P,V)$ satisfies $f(T_{0})=g(T_{0})$. Hence, $f_{\mathfrak{m}}$ is onto if $\mathfrak{m}$ is not among $\mathfrak{m}_{1},\dots,\mathfrak{m}_{k}$. Moreover, $f(T_{i})=(g+\sum_{j=1}^{i}f_{j})(T_{i})$ shows that $f_{\mathfrak{m}_{i}}$ is onto for every $i=1,\dots,k$. This implies that $f$ is onto.
Assume $f_{1},\dots,f_{\ell-1}$ and $T_{0},T_{1},\dots,T_{\ell-1}$ are found. Consider $g_{\ell}=g+\sum_{i=1}^{\ell-1}f_{i}$. Recall that $V_{\mathfrak{m}_{\ell}}$ is a finitely generated projective module over a semiperfect ring $\Lambda_{\mathfrak{m}_{\ell}}$. Then $V_{\mathfrak{m}_{\ell}}/V_{\mathfrak{m}_{\ell}}J(\Lambda_{\mathfrak{m}_{\ell}})$ is a finitely generated (semisimple module). Therefore, if $\pi\colon V_{\mathfrak{m}_{\ell}}\to V_{\mathfrak{m}_{\ell}}/V_{\mathfrak{m}_{\ell}}J(\Lambda_{\mathfrak{m}_{\ell}})$ is the canonical projection, $\pi((g_{\ell}(P))_{\mathfrak{m}_{\ell}})$ is a finitely generated submodule of $V_{\mathfrak{m}_{\ell}}/V_{\mathfrak{m}_{\ell}}J(\Lambda_{\mathfrak{m}_{\ell}})$. Let $T_{\ell}^{\prime}\subseteq P$ be a finitely generated submodule of $P$ containing $T_{\ell-1}$ such that $(g_{\ell}(T_{\ell}^{\prime}))_{\mathfrak{m}_{\ell}}+V_{\mathfrak{m}_{\ell}}J(\Lambda_{\mathfrak{m}_{\ell}})=(g_{\ell}(P))_{\mathfrak{m}_{\ell}}+V_{\mathfrak{m}_{\ell}}J(\Lambda_{\mathfrak{m}_{\ell}})$.
By Lemma 6.13, there exists $f_{\ell}\in{\rm Hom}_{\Lambda}(P,V)$ such that $(f_{\ell})_{\mathfrak{m}_{\ell}}$ is onto and $f_{\ell}(T_{\ell}^{\prime})=0$. Then, the image of $g_{\ell}+f_{\ell}$ contains $g_{\ell}(T_{\ell}^{\prime})$. Hence, $\mbox{\rm Im}\ (g_{\ell}+f_{\ell})_{\mathfrak{m}_{\ell}}+V_{\mathfrak{m}_{\ell}}J(\Lambda_{\mathfrak{m}_{\ell}})$ also contains $\mbox{\rm Im}\ (g_{\ell})_{\mathfrak{m}_{\ell}}$, and since $(f_{\ell})_{\mathfrak{m}_{\ell}}$ is onto, $\mbox{\rm Im}\ (g_{\ell}+f_{\ell})_{\mathfrak{m}_{\ell}}+V_{\mathfrak{m}_{\ell}}J(\Lambda_{\mathfrak{m}_{\ell}})=V_{\mathfrak{m}_{\ell}}$. Nakayama’s Lemma gives that $(g_{\ell}+f_{\ell})_{\mathfrak{m}_{\ell}}$ is onto. Therefore, there exists a finitely generated submodule $T_{\ell}$ of $P$ containing $T_{\ell}^{\prime}$ such that $((g_{\ell}+f_{\ell}))_{\mathfrak{m}_{\ell}}((T_{\ell})_{\mathfrak{m}_{\ell}})=V_{\mathfrak{m}_{\ell}}$, that is, $(iii)$ holds if $i=\ell$.
@@
Proof. Propositions 6.12 and 6.16 show that $V_{r}^{*}(\Lambda)=V_{r}(\Lambda)\sqcup B_{r}(\Lambda)$. The symmetry of the hypothesis implies that $V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)=V_{\ell}^{*}(\Lambda)$. By Remark 6.10, we can deduce from Corollary 2.9 that both monoids are isomorphic.
Hinohara [hinohara] proved that every projective module over an $h$-local (or, more generally, weakly noetherian) domain is either finitely generated or free. We can extend this result to our setting.
@@
[hinohara] Let $R$ be an $h$-local domain with field of fractions $Q$, and let $\Lambda$ be a locally semiperfect torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian. For each maximal ideal $\mathfrak{m}$ of $R$, assume that $\Lambda_{\mathfrak{m}}$ is Morita equivalent to a local ring. Then any infinitely generated projective $\Lambda$-module is free.
Proof. By Remark 6.10, ${\mathcal{T}}(\Lambda)=\{0,\Lambda\}$, and hence $V(\Lambda)\sqcup B(\Lambda)=V(\Lambda)\sqcup\{[\Lambda^{(\omega)}]\}$. The result then follows from Theorem 6.17.